OpenStudy (anonymous):

find the solution set for x, y, z 2x + y - 2z= 3 3x - y - 3z= 2 5x + 6y -2z= 2

OpenStudy (snowfire):

Do you know how to use augmented matrices?

OpenStudy (wolf1728):

You will have to solve for x, y and z separately. You would start be eliminating 1 unknown from the equations: A) 2x + y - 2z= 3 B) 3x - y - 3z= 2 C) 5x + 6y -2z= 2 Multiply equation C by minus 1 then add all 3 equations: A) 2x + y - 2z= 3 B) 3x - y - 3z= 2 C) -5x - 6y +2z= -2 Total of all 3 = -6y -3z = 3 -3z = 6y +3 -z = 2y + 1 or z = -2y -1 inputting this value of z into equation A and B A) 2x + y - 2z= 3 B) 3x - y - 3z= 2 2x + y + 4y +2 = 3 3x -y + 6y + 3 = 2 Multiplying equation B by -2/3 -2x + (2/3)y -4y -2 = -4/3 then adding equation A 2x + y + 4y +2 = 3 sums to (5/3)y = (5/3) Y = 1 We have solved the first unknown

OpenStudy (wolf1728):

Can you solve for x and z?

OpenStudy (snowfire):

Oh wow, you didn't hold back wolf man

OpenStudy (wolf1728):

Thanks snowfire

OpenStudy (e.mccormick):

@Snowfire \(\left[\begin{array}{ccc|c} 2 & 1 &- 2& 3 \\ 3 & -1& - 3& 2 \\ 5 & 6 & -2 & 2 \end{array}\right] \) Now they just need to put it in rref. \(\large\ddot \smile\)